OpenStudy (anonymous):

Verify implicit solution dX/dt=(x-1)(2x-1), ln(2x-1/x-1)=t

OpenStudy (anonymous):

\[\ln(\frac{2x-1}{x-1})=\ln(2x-1)-\ln(x-1)\]

OpenStudy (anonymous):

t=ln(2x-1)-ln(x-1)

OpenStudy (anonymous):

\[\frac{d}{dt}[t]=\frac{d}{dt}[\ln(2x-1)-\ln(x-1)]\]

OpenStudy (ash2326):

Yeah you're right:)

OpenStudy (anonymous):

i'm getting a negative somewhere just writing it out

OpenStudy (anonymous):

\[1=\frac{2}{2x-1}\frac{dX}{dt}-\frac{1}{x-1}\frac{dX}{dt}\]

OpenStudy (anonymous):

\[1=(\frac{2}{2x-1}-\frac{1}{x-1})\frac{dX}{dt}\]

OpenStudy (anonymous):

\[1=\frac{2(x-1)-1(2x-1)}{(2x-1)(x-1)}\]

OpenStudy (anonymous):

\[1=\frac{2x-2-2x+1}{(2x-1)(x-1)}\]

OpenStudy (anonymous):

\[1=\frac{-1}{(2x-1)(x-1)}\]

OpenStudy (anonymous):

dX/dt of course

OpenStudy (anonymous):

\[1=\frac{-1}{(2x-1)(x-1)}\frac{dX}{dt}\]

OpenStudy (anonymous):

\[-(2x-1)(x-1)=\frac{dX}{dt}\] ? was everything done right? is this an error in the book?

OpenStudy (ash2326):

Yeah you have done everything correctly

OpenStudy (anonymous):

so it's an error in the book? as -(2x-1)(x-1) does not equal (2x-1)(x-1)

OpenStudy (ash2326):

Yeah

OpenStudy (anonymous):

i'll have my teacher know lol